# You Only Click Once: Single Point Weakly Supervised 3D Instance Segmentation for Autonomous Driving

### v2

@@ 6.2 More Experiments
Results on nuScenes. We conduct experiments on the nuScenes dataset, using only the point clouds projected onto the image. Data augmentation for single-frame data included global rotation, translation, scaling, random flipping, and shuffling. As shown in Table 8, our method YoCo achieves a 14% improvement in mAP and a 17.868% improvement in mIoU compared to the Click-based method, demonstrating its effectiveness.
Sup. Method mAP mIoU Full SparseUnet 49.56 68.056 Cylinder3D 45.74 66.198 PTv3 46.1 71.972 Weak Click 19.35 37.686 YoCo (Ours) 33.35 55.554
Table 8: Performance Comparisons on NuScenes Validation Dataset.
Greater click range error. To evaluate the robustness of YoCo for larger target sizes in autonomous driving, we conduct additional experiments with error ranges of 1m, 1.5m, and 2m for vehicle classes, as shown in Table 9. Expanding the click range to 2.0m results in stable performance and even improvement, demonstrating the robustness of our method. These experiments suggest that the BEV center is not the optimal click, and identifying the optimal click will be explored in future work.
@@
Discard click analysis. To verify whether all foreground objects need to be clicked, we conduct a random discard experiment. It is noteworthy that our default method inherently discards objects with insufficient point cloud representations in the Bird’s Eye View (BEV) perspective, resulting in a 16.31% object discard rate. We further evaluate the impact of additional discarding by randomly removing instances with probabilities of 10%, 20%, and 30%, as detailed in Table 11 (b), (c), and (d). The experimental results indicate that as the discard rate increases, the network performance gradually declines. However, even with a discard proportion of 37.19%, our method still achieves 53.96% mAP and 73.248% mIoU, demonstrating the robustness of YoCo.
@@
Table 13: Performance Comparison of Supervision Strategies on the Waymo Validation Dataset. indicates pseudo labels generated by SAM using the corresponding click annotations as prompts, represents pseudo labels produced by our proposed VFM-PLG module, and YOLO refers to pseudo labels derived from YOLO prediction results.
@@ 6.3 Visualization
In Figure 6, we evaluate the performance of SparseUnet trained with various annotation strategies. Our proposed method exhibits superior visual quality, significantly outperforming click-based approaches by generating more accurate and spatially coherent predictions. This improvement is largely driven by the integration of deep prior information into the pseudo label generation process, which enhances the quality of the generated labels.

### v3

@@ 6.2 More Experiments
Results on nuScenes. We conduct experiments on the nuScenes dataset, using only the point clouds projected onto the image. Data augmentation for single-frame data included global rotation, translation, scaling, random flipping, and shuffling. We specifically focus on the barrier, bicycle, bus, car, construction vehicle, motorcycle, pedestrian, traffic cone, trailer, and truck categories for experiments. As shown in Table 8, our method YoCo achieves a 14% improvement in mAP and a 17.868% improvement in mIoU compared to the Click-based method, demonstrating its effectiveness.
Sup. Method mAP mIoU mIoU (fg) Full SparseUnet 49.56 68.056 78.548 Cylinder3D 45.74 66.198 76.335 PTv3 46.1 71.972 81.264 Weak Click 19.35 37.686 58.067 YoCo (Ours) 33.35 55.554 68.267
Table 8: Performance Comparisons on NuScenes Validation Dataset. fg denotes that only foreground points participate in the evaluation, ignoring background points.
Greater click range error. To evaluate the robustness of YoCo for larger target sizes in autonomous driving, we conduct additional experiments with error ranges of 1.0m, 1.5m, and 2.0m for vehicle classes, as shown in Table 9. Expanding the click range to 2.0m results in stable performance and even improvement, demonstrating the robustness of our method. These experiments suggest that the BEV center is not the optimal click, and identifying the optimal click will be explored in future work.
@@
Discard click analysis. To verify whether all foreground objects need to be clicked, we conduct a random discard experiment. It is noteworthy that our default method inherently discards objects with insufficient point cloud representations in the Bird’s Eye View (BEV) perspective, resulting in a 16.31% object discard rate. We further evaluate the impact of additional discarding by randomly removing instances with probabilities of 10%, 25%, and 30%, as detailed in Table 11 (b), (c), and (d). The experimental results indicate that as the discard rate increases, the network performance gradually declines. However, even with a discard proportion of 37.19%, our method still achieves 53.96% mAP and 73.248% mIoU, demonstrating the robustness of YoCo.
@@
Table 13: Performance Comparison of Supervision Strategies on the Waymo Validation Dataset. YOLO refers to pseudo labels derived from YOLO prediction results.
@@ 6.3 Visualization
In Figure 6, we evaluate the performance of SparseUnet trained with various annotation strategies. Our proposed method exhibits superior visual quality, significantly outperforming click-based approaches by generating more accurate predictions. This improvement is largely driven by the integration of deep prior information into the pseudo label generation process, which enhances the quality of the generated labels.